OpenStudy (anonymous):

Solve: log (x + 6) – log (x + 2) = log x

OpenStudy (ankit042):

log(a) - log(b) = log(a/b) use this and then cancel logs on both sides

OpenStudy (anonymous):

very nice, i was confused how to get rid of logx

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i know that minus = divide but how to get rid of log on the =logx

OpenStudy (ankit042):

log(A) = log(b) A = b

OpenStudy (anonymous):



OpenStudy (anonymous):

so the log just cancels?

OpenStudy (anonymous):

(+6)/(x+2)=x

OpenStudy (ankit042):

yes x+6/x+2 = x

OpenStudy (anonymous):

(x=6)/(x+2)=x

OpenStudy (anonymous):

tyty

OpenStudy (ankit042):

npnp :D